OpenStudy (anonymous):

Which of the following would not be the result of a substitution in the following system? 2x + y = 7 y - x = 1 7 - 2x - x = 1 2x + x + 1 = 7 2(y + 1) + y = 7

OpenStudy (anonymous):

its not 1

OpenStudy (anonymous):

why? 2x+y=7 y-x=1 y=1+x 2x+x+1=7

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

please really excuse when i do mistakes)

jigglypuff314 (jigglypuff314):

@hhelpplzzzz @AlexKross still incorrect, please read the question carefully: "following would NOT be the result "

OpenStudy (anonymous):

okay thanks